Selection Sort is a common algorithm to sort the elements of an array. First, it scans the elements of the array to find the smallest value and places it at index 0, thus creating a sorted “subarray” of size 1 that contains the smallest value. Then it scans the remaining unsorted values for the new smallest value and places it at index 1, creating a sorted subarray of size 2 that contains the 2 smallest values. It continues in this way until the sorted subarray grows to contain all of the array elements, and therefore the entire array is sorted. Based on this description, what is the best big O estimate for Selection Sort? (Try to think about this for a bit before looking up the solution)

O(n)

O(log n)

O(n log n)

O(n2)
Answer #1


d)

O(n2)

each iteration for finding smallest number taken O(n) complexity.

since this is done n times.

total time complexity = n*O(n) = O(n2)